inspiratory time 1.1 s, tidal volume 0.9 L inspiratory flow rate L/min inspiratory time 0.4 s, tidal volume 200 mL inspiratory flow rate L/min inspiratory time 0.5 s, tidal volume 0.8 L inspiratory flow rate L/min inspiratory time 1.5 s, tidal volume 925 mL inspiratory flow rate te/min Tidal volume 0.9 L, flow rate 45 L/min Inspiratory time is seconds Tidal volume 0.8 L, flow rate 60 L/min Inspiratory time is seconds Tidal volume 700 l-mL, flow rate 42 L/min Inspiratory time iss seconds
Answer #1

ANSWER

    INSPIRATORY FLOW RATE.

               Formula

                   Inspiratory flow rate = Tidal volume (L) / Inspiratory time (s) * 60

As per formula,

       * Inspiratory time = 1.1 s. , Tidal volume = 0.9 L.

          Inspiratory flow rate = 0.9 / 1.1 * 60

                                        = 0.818 * 60

                                           = 49.08 L/mt.

      * Inspiratory time = 0.4 s, Tidal volume = 200 ml (0.2 L).

                 Inspiratory flow rate = 0.2 / 0.4 * 60

                                                 = 0.5 * 60

                                                = 30 L/mt.

     * Inspiratory time = 0.5 s ,Tidal volume = 0.8 L.

                Inspiratory flow rate = 0.8 / 0.5 * 60

                                                  = 1.6 * 60

                                                = 96 L /mt.

    * Inspiratory time = 1.5 s ,Tidal volume = 925 ml ( 0.925ml).

                      Inspiratory flow rate = 0.925 / 1.5 * 60

                                                        = 0.617 * 60

                                                        = 37.02 L /mt.

* INSPIRATORY TIME.

             Formula

                   Inspiratory time = Tidal volume (L) / Flow rate (L/mt) * 60

      * Tidal volume = 0.9 L , Flow rate = 45 L / mt.

                 Inspiratory time = 0.9 / 45 * 60

                                             = 0.02 * 60

                                             = 1.2 sec.

     * Tidal volume = 0.8 L, Flow rate = 60 L /mt..

               Inspiratory time = 0.8 / 60 * 60

                                       = 0.013 * 60

                                         = 0.78 sec.

     * Tidal volume = 700 ml (0.7 L) , Flow rate = 42 L /mt.

               Inspiratory time = 0.7 / 42 * 60

                                          = 0.017 * 60

                                        = 1.02 sec.